reduce x1-2y+3z-4x2x3+6x1x3 to canonical forms


1
Expert's answer
2021-06-16T04:53:35-0400

Reduce the quadratic form "Q(x,y,z)=x^2+5y^2+z^2+2xy+2yz+6xz" to canonical form.

The given quadratic form is


"Q(x, y, z)=x^2+5y^2+z^2+2xy+2yz+6xz"

The matrix of the given quadratic form is

"A=\begin{pmatrix}
 1 & 1 & 3 \\
 1 & 5 & 1\\
 3 & 1 & 1 \\
\end{pmatrix}"

"A-\lambda I=\begin{pmatrix}
 1-\lambda & 1 & 3 \\
 1 & 5-\lambda & 1\\
 3 & 1 & 1-\lambda \\
\end{pmatrix}"

"\det(A-\lambda I)=\begin{vmatrix}
 1-\lambda & 1 & 3 \\
 1 & 5-\lambda & 1\\
 3 & 1 & 1-\lambda \\
\end{vmatrix}"

"=(1-\lambda)\begin{vmatrix}
 5-\lambda& 1 \\
 1 & 1-\lambda
\end{vmatrix}-\begin{vmatrix}
 1 & 1 \\
 3 & 1-\lambda
\end{vmatrix}+3\begin{vmatrix}
 1 & 5-\lambda \\
 3 & 1
\end{vmatrix}"

"=(1-\lambda)(5-5\lambda-\lambda+\lambda^2-1)-(1-\lambda-3)"


"+3(1-15+3\lambda)=4-6\lambda+\lambda^2-4\lambda+6\lambda^2-\lambda^3"

"+2+\lambda-42+9\lambda"

"=-\lambda^3+7\lambda^2-36=0"


"-\lambda^2(\lambda-6)+(\lambda-6)(\lambda+6)=0"

"-(\lambda-6)(\lambda^2-\lambda-6)=0"

"-(\lambda-6)\lambda-3)(\lambda+2)=0"

"\lambda_1=6, \lambda=3, \lambda_3=-2"

These are the eigenvalues.

"\lambda=6"


"\begin{pmatrix}
 1-\lambda & 1 & 3 \\
 1 & 5-\lambda & 1\\
 3 & 1 & 1-\lambda \\
\end{pmatrix}=\begin{pmatrix}
 -5 & 1 & 3 \\
 1 & -1 & 1\\
 3 & 1 & -5 \\
\end{pmatrix}"

"\begin{pmatrix}
 -5 & 1 & 3 \\
 1 & -1 & 1\\
 3 & 1 & -5 \\
\end{pmatrix}\to\begin{pmatrix}
 1 & 0 & -1 \\
 0 &1 & -2\\
 0 & 0 & 0 \\
\end{pmatrix}"

The eigenvector is "\vec v=\begin{pmatrix}
 1 \\
 2 \\
1
\end{pmatrix}"


"\lambda=3"


"\begin{pmatrix}
 1-\lambda & 1 & 3 \\
 1 & 5-\lambda & 1\\
 3 & 1 & 1-\lambda \\
\end{pmatrix}=\begin{pmatrix}
 -2 & 1 & 3 \\
 1 & 2 & 1\\
 3 & 1 & -2 \\
\end{pmatrix}"

"\begin{pmatrix}
 -2 & 1 & 3 \\
 1 & 2 & 1\\
 3 & 1 & -2 \\
\end{pmatrix}\to\begin{pmatrix}
 1 & 0 & -1 \\
 0 &1 & 1\\
 0 & 0 & 0 \\
\end{pmatrix}"

The eigenvector is "\vec u=\begin{pmatrix}
 1 \\
 -1 \\
 1
\end{pmatrix}"


"\lambda=-2"


"\begin{pmatrix}
 1-\lambda & 1 & 3 \\
 1 & 5-\lambda & 1\\
 3 & 1 & 1-\lambda \\
\end{pmatrix}=\begin{pmatrix}
 3 & 1 & 3 \\
 1 & 7 & 1\\
 3 & 1 & 3 \\
\end{pmatrix}"

"\begin{pmatrix}
 3 & 1 & 3 \\
 1 & 7 & 1\\
 3 & 1 & 3 \\
\end{pmatrix}\to\begin{pmatrix}
 1 & 0 & 1 \\
 0 &1 & 0\\
 0 & 0 & 0 \\
\end{pmatrix}"

The eigenvector is "\vec w=\begin{pmatrix}
 -1 \\
 0 \\
 1
\end{pmatrix}"

Form the matrix "P," whose column "i" is eigenvector no. "i"


"P=\begin{pmatrix}
 1 & 1 & -1 \\
 2 & -1 & 0 \\
 1 & 1 & 1 \\
\end{pmatrix}"

Form the diagonal matrix "D," whose element at row "i," column "i" is eigenvalue no. "i"


"D=\begin{pmatrix}
 6 & 0 & 0 \\
 0 & 3 & 0 \\
 0 & 0 & -2 \\
\end{pmatrix}"

The matrices "P" and "D" are such that


"P^{-1}AP=D"

Hence the quadratic form is reduced to a sum of squeres, i.e. canonical form:


"6\tilde{x}^2+3\tilde{y}^2-2\tilde{z}^2"

"P=\begin{pmatrix}
 1 & 1 & -1 \\
 2 & -1 & 0 \\
 1 & 1 & 1 \\
\end{pmatrix}" is the matrix of transformation which is an orthogonal matrix.

The canonical form of the quadratic form is


"6\tilde{x}^2+3\tilde{y}^2-2\tilde{z}^2"